the scale is 2 in : 3 ft. perimeter 51 in.

Area: 184 in^2
the builder also needs to know how much carpet to put in the room. what is the actual area of the room to the nearest square foot?

Answer by Theo(13342)

the scale is 2 inches on the diagram equals 3 feet in the real world.

this give you 1.5 feet per inch.

51 inches * 1.5 = 76.5 feet

the perimeter is 76.5 feet.

the area is 184 square inches.

1 inch is equal to 1 inch * 1 inch = 1 square inch

1 inch is equal to 1.5 feet, therefore 1 square inch = 1.5 * 1.5 = 2.25 square feet.

therefore 184 square inches is equal to 2.25 * 184 = 414 square feet.

you could also do it another way.

184 is equal to 2 * 92

since 1 inch = 1.5 feet, then you get;

2 inches * 92 inches = 3 feet * 138 feet = 414 square feet.

it's just another way to get the same answer.